40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog, Page 3 of 38  Not logged in ELOG logo
ID Date Author Type Category Subject
  1869   Tue Nov 24 16:30:59 2020 PacoMiscEquipmentLoanbroadband EOM

Borrowed 1 (new focus) broadband EOM from CTN for temporary use in Crackle (2 um OPO exp)

  1868   Tue Nov 24 12:32:56 2020 KojiMiscGeneralMDT694B python controller

Wow! This is really cool! I didn't realize that this small box has such many remote capabilities.
We have this piezo controller everywhere in the labs and your code gives us a lot of opportunities to implement process automation.

  1867   Mon Nov 23 11:44:13 2020 PacoMiscGeneralMDT694B python controller

Today entered lab ~ 09:00. Over the weekend I coded a PySerial wrapper for the thorlabs MDT694B single channel piezo controller. I spent some time testing and debugging the code but it now works fine (tested on Linux, python=3.8.6 and PySerial=3.4-4). The wrapper refers to the manual available here. The code is available in the labutils repo

  1866   Mon Nov 23 11:43:26 2020 PacoMiscEquipmentLoanbeam profiler

Shruti took back the beam profilers today AM to Cryo.

 


Shruti: returned to Gabriele's office

Quote:

I placed the two beam profilers with the two laptops and chargers right inside the Crackle lab, as requested by Paco.

 

  1865   Fri Nov 20 18:43:31 2020 KojiSummaryGeneralDelivered the instrument / components for 2um ECDL

> Temp Controller / TED200C / Qty 2 Note One unit temporary used by 2um PD test setup

I brought the brand new TED200C from QIL to Crackle (Permanent move).

The unit used for 2um PD test setup will stay in QIL (Permanent)

 

  1864   Wed Nov 18 17:49:05 2020 PacoDailyProgressOpticsbeam profiles

Used BeamR and WinCamD to profile the two light sources (ECDL and OPO pump) 

(1) ECDL; profile 19** nm beam after the aspheric lens. I guess we want this beam to be nominally collimated for optimal feedback with the Littrow-configured grating, so I recorded the 1/e^2 waists (x, y) as a function of longitudinal displacement. The result is attached below. Linear fits provide rough estimates for the beam divergences, giving 2.0 mrad along x (parallel to the table) and 1.2 mrad along y (normal to the table) suggesting some astigmatism which is common in high NA aspheric lenses. I inspected the distance from the aspheric lens to the SAF gain chip and measured ~ 2.0 mm (compared to the 1.99 mm working distance specified for this lens with NA=0.71). The SAF1900S specifies a beam divergence angle of 35 deg  (corresponding to NA=0.57), so there is room for improvement by tweaking the aspheric flexure alignment.

(2) NPRO; profile 1064 nm beam at low power (~10 mW) right after the head output. Having 10x more power made things way easier for this as compared to the ECDL, but the method was the same (record 1/e^2 waists as a function of longitudinal displacement). The result is attached below. Linear fits provide rough estimates for the beam divergences, giving 2.1 mrad along x, and 2.1 mrad along y. Here I grabbed the specified divergence of 2.3 mrad from a relatively old manual, and even drew the displaced waist profile (w0 = 160 um) which seemingly fit the profile, but the actual values may be different.

Attachment 1: 2020_11_18_ecdlprofile.png
2020_11_18_ecdlprofile.png
Attachment 2: 2020_11_18_mephistoprofile.png
2020_11_18_mephistoprofile.png
  1863   Wed Nov 18 14:29:00 2020 KojiMiscEquipmentLoanbeam profiler

Note: Please don't try to connect these old Windows to the network. We just extract the data via USB etc, and that's all the connection we allow.

  1862   Wed Nov 18 09:14:11 2020 shrutiMiscEquipmentLoanbeam profiler

I placed the two beam profilers with the two laptops and chargers right inside the Crackle lab, as requested by Paco.

  1861   Wed Nov 11 12:55:18 2020 PacoLab InfrastructureGeneral1064 pump power curve

Log of the output power vs current in the 1064 nm (Innolight) pump laser. The crystal temperature was set to 45.5 C, and the current limit is set to 2.1 A

Attachment 1: 2020_11_10_mephisto_test.png
2020_11_10_mephisto_test.png
  1860   Wed Nov 11 12:28:33 2020 PacoDailyProgressGeneralEP30 softening

Attempted two methods to soften EP30-2, the results are summarized below.

(a) Heat -- Used the heat gun set to 200 F (~ 93 C) and held it near the back of the part so that the grating surface was never in direct exposure. The airflow was kept constant for a period of ~ 10 min, while I periodically checked to see if there were any signs of bond softening. After no signs of softening, I stopped and moved to method (b).

(b) Solvent -- After brief investigation and referring to T1400711, I got some acetone from CTN, and set up a ~ 50 ml bath. The part was not completely submerged and was arranged such that the grating face was always exposed to air, which I left for ~14 hours. The drawback of this method is that some of the acetone evaporated and at some point the EP30 bond stopped being in contact with the solvent. A picture for reference is attached, with the light blue line indicating the highest acetone level, and the red line indicating the EP30 bond level at the beginning of the bath.

Attachment 1: acetone_bath.jpg
acetone_bath.jpg
  1859   Tue Nov 10 10:32:57 2020 PacoDailyProgressGeneralSAF gain chip first light

Set grating in front of 1064 nm beam (current set to 1.058 A for a beam visible on the IR card). After testing both orientations, it becomes clear the grating is misoriented. The difference is very clear, there is only specular reflection in the current configuration, whereas the m=0, and +- 1 orders are visible in the 180 deg flipped configuration. 

Quote:

Very exciting to see the gain chip curve!

Grating orientation: Whaaat... If you already have the 1um laser SOP approved, you can use that laser to check the grating orientation.

 

 

  1858   Tue Nov 10 00:02:20 2020 KojiDailyProgressGeneralSAF gain chip first light

Very exciting to see the gain chip curve!

Grating orientation: Whaaat... If you already have the 1um laser SOP approved, you can use that laser to check the grating orientation.

 

  1857   Mon Nov 9 17:38:18 2020 PacoDailyProgressGeneralSAF gain chip first light

Entered Crackle ~ 8:47 AM.

Briefly fixed the LDC220C connection to the SAF1900 as described previously, and then installed the aspheric flexture and shoulder to the assembly (pictures below). Then, I used the thermal power meter head borrowed from OMC to check for emission as a function of laser diode current at a fixed temperature of 25 C (to match testing conditions). The result is below, where I seem to be getting slightly better amplified spontaneous emission (ASE) power than the attached test sheet. It may as well be that I am not measuring the ASE power alone, but I cannot presently determine this.

I added the grating and moved the power meter to the correct output aperture, but failed to detect any power. This suggests a wrong grating orientation, although I will try to verify this more carefully.

Attachment 1: 2020_11_09_ldc200c_test.png
2020_11_09_ldc200c_test.png
Attachment 2: 20201109_aspheric.jpg
20201109_aspheric.jpg
Attachment 3: 20201109_firstlight.jpg
20201109_firstlight.jpg
Attachment 4: 20201109_grating.jpg
20201109_grating.jpg
  1856   Fri Nov 6 11:58:02 2020 KojiLab InfrastructureGeneralECDL Gratings / Lenses delivered. Powermeter lending from OMC Lab

The gratings and aspheric lenses glued on the mounts were delivered to the lab on Thu.

The powermeter controler + S401C head was lent from OMC Lab. Returned to OMC Jul 15, 2020 KA

Attachment 1: P_20201105_195258.jpg
P_20201105_195258.jpg
  1855   Fri Nov 6 10:47:49 2020 PacoDailyProgressGeneralConnecting SAF1900S

Here is a summary for how to connect the SAF1900S gain chip to TED200C temperature controller and LDC220C diode current driver. The chip itself lacked substantial documentation, so this comes after requesting tech support from the manufacturer. The SAF1900S pinout is

1 - TEC+ 

6 - TEC-

2 - Thermistor

3 - Thremistor

4 - Anode

5 - Cathode

The TED200C has a DSub15 output, but the cable provides a DSUB9 adapter. Then, only the following pins are connected to the SAF1900S

2: Thermistor

3: Thermistor

4: TEC+

5: TEC-

The LDC220C has a DSUB9 output, and its bipolar nature allows it to drive either anode-grounded or cathode-grounded diodes, so the question was wether the SAF1900S is AG/CG? In a first attempt, I assumed the diode was meant to be driven with a floating source (and that the LDC220C could do that), but the driver remained in "LD OPEN" state. Then, I revised the documentation for TLK1900 (an old, discontinued laser kit using the same gain chips). There, the bottom line seemed to suggest CG, but to be sure I asked a technician in thorlabs. They say most of their 14 pin butterfly chips are AG, but the 6pin ones seem to be CG. Anyways, the relevant pins (for either connection) are:

3: Ground (for AG/CG)

7: LD Cathode (for floating / AG)

8: LD Anode (for floating / CG)

After some communication with ANU's Disha, I found the diode pins are floating from the case (personally confirmed this), and an additional connection between pins 1 and 5 of the LDC220C needs to be established to override the interlock. The suggested connections are three: shortcut, resistance < 430 Ohm, or LED || 1 kOhm resistor (to match the Laser ON status in the front panel). I opted for this last one, made the connections and was able to correctly feed the SAF gain chip.

Attachment 1: saf1900s_conn.png
saf1900s_conn.png
Attachment 2: saf1900_conn2.png
saf1900_conn2.png
  1854   Wed Nov 4 01:25:37 2020 KojiSummaryOpticsAspheric Lenses and Grating Plates bonded on their mounts

Attachment 1: Black Diamond (GeSbSe) Lens was mounted on the flexure mount. The flat surface should face to the gain chip. It was aligned on the wipe to be flush with the protrusion.

Attachment 2: Applied glue on the four grooves of each flexure mount.

Attachment 3: The grating was bonded on the mount. The arrow marks were arranged as Paco directed. The mount could not stand by itself. And the screws were placed to stop the grating skating on the mount.

Attachment 1: IMG_0188.jpg
IMG_0188.jpg
Attachment 2: IMG_0208.jpg
IMG_0208.jpg
Attachment 3: IMG_0202.jpg
IMG_0202.jpg
  1853   Mon Nov 2 18:23:46 2020 PacoDailyProgressGeneral 

Today; entered lab at ~09:08. I verified the orientations of the aspheric lenses and blaze gratings relative to the flextures, packaged and then dropped the parts for epoxying to Koji in 40m ~ 11:00. Spent some time between 12:00 and 12:45 finishing the ECDL connections. Everything looked good so I hooked it up to the TED200C controller. After a bit of research, I found out the Steinhart constants for the 10k thermistor;

a=1.129241e-3

b=2.341077e-4

c=8.775468e-8

Plugging these into the Steinhart equation give the actual temperatures from the Tact output on the TED200C (otherwise read as kOhm). According to the spec sheet, the TEC was tested at 250 mA (0.40 V), so not knowing a bunch more, set I_TEC on the TED200C to this limit and inspect the actual TEC current by scanning the Tset (setpoint) and recording the current in the ~ 15 - 25 deg C (attached plot, horizontal line marks room temperature). The diode current driver is hooked up, and everything is on the table as is. Left Crackle ~ 18:30.

Attachment 1: 2020_11_02_ted200c_test.png
2020_11_02_ted200c_test.png
  1852   Wed Oct 28 17:29:14 2020 PacoSummaryGeneralLab organization

I entered Crackle lab circa ~11:15. I started some basic lab inventory and started cleaning/organizing stuff. We will use the first optical table (as you enter the lab) because it's the easiest to clear (see below before and after clearing). Some of the cleared items on the table include:

- UHV foil (moved to top left cabinet above the work bench)
- OSEM components for Crackle (?) (moved to top left cabinet above the work bench)
- Various metallic parts/components (moved some in a plastic container to the second drawer from the bottom of the second red tool storage, and others to the second optical table)
- Various screw/screwdriver kits (moved some to work bench right by the electronic storage area and others to the second optical table)
- Power supply and laser diode driver (moved to control/acquisition rack)

I then moved the 1064nm pump Innolight Mephisto 800NE to the clear table, clamped it down, and cleaned/organized the lab a little, which involved:

- Shelve orphan/incomplete PCBs and electronic components from the work bench up to the cabinets.
- Organized some cables by the fume hood.
- Organized other random hardware on the work benches. 

I found the Emergency STOP (OMRON STI #A22EM02B) button buried on the fume hood, so I gave it a quick test, and after confirming it worked I wired it to the interlock on the back of the laser controller. Then tested it along with the interlock and verified it's working, but I have yet to solder it nicely (I didn't commit to the wire lengths yet).

Left at ~ 14:45. Noted that we had more cockroaches in the floor at the beginning of the day than 2 um laser sources. Now we are tied.

Attachment 1: IMG-20201028-WA0011.jpeg
IMG-20201028-WA0011.jpeg
Attachment 2: IMG_20201028_144653.jpg
IMG_20201028_144653.jpg
  1851   Tue Oct 20 20:41:59 2020 KojiSummaryGeneralDelivered the instrument / components for 2um ECDL

I brought in the instrument and components for 2um ECDL:

1. SAF Gain chip / SAF1900S / Qty2
2. Grating / GR25-0616 / Qty2
3. 3axis piezo mount / POLARIS-K1S3P / Qty2
4. Lens / 390093-D / Qty2
5/6 Thorlabs small components / F3ES20, F3ESN1P / Qty2 ea
7 N/A
8~13 Machines Metal Components / D1900435, D1900429, D1900433, D1900432, D1900430, D1900434 / Qty 2ea
14~17 McMaster Carr fastners / 92196a192, 92196a110, 92196a079, 92196a081 / Qty 100 ea
18 Temp Controller / TED200C / Qty 2 Note One unit temporary used by 2um PD test setup
19 Laser Current Driver / LDC220C / Qty 2
20 Piezo Driver MDT694B / Qty 2

Attachment 1: 20201020192651_IMG_0165.JPG
20201020192651_IMG_0165.JPG
  1850   Mon Apr 20 22:56:30 2020 ranaDailyProgressNoise BudgetNoisemon:DAC noise analysis from L1 and H1

for some reason the DAC noise estimate is too high, it can't really be so large compared to the real DARM curve (see the noise budget curves from LLO - there are other noise sources besides DAC noise)

some possibilities:

  1. maybe the DAC noise calibration into meters is wrong? I can't tell from the code where this came from. It would be good to put a comment in there.
  2. perhaps most of this noise is actually angular noise. The ASC control signals are adjusted by tuning the digital coefficients (before MASTER_OUT) so that the angle to length coupling is minimized. I think something like this has to be done to remove the angular noise from the DAC noise estimate.
  3. internal saturation of the DAC noise monitor?

I hve modified the code to plot nicer and also to remove some divide by zero problems. There is also still some warnings about other divide by zero - those should probably be fixed by examining how better to handle it when the coherence goes to zero.

Attachment 1: DACnoise-comparison.pdf
DACnoise-comparison.pdf
Attachment 2: dacNB.zip
  1849   Sat Apr 18 16:21:32 2020 DuoComputingNoise BudgetNoisemon:DAC noise analysis from L1 and H1

There we go. Based on the noisemon data at L1 and H1, I calculated the DAC noises at those sites, using roughly the same approach as described in 1847.

I used the coherence between the master channel and the noisemon channel to calculate the total noise going into the coils.

Then I converted the ADC noise and noisemon noise to DAC volts and subtracted them from the total noise. I compared the result of the subtraction, which should be DAC noise, at least in the passband (20-100Hz), with the G1401399 model and made a noise budget, shown in attachment 1. We can see that, as designed, the DAC noise is sufficiently amplified so that it dominates over the noisemon noise or the ADC noise in the passband.

Next, I projected the DAC noise to strain noise and summed them up for all the four channels in all the four stations.

Finally, I compared this with the interferometer noise spectrum based on data in L1:OAF-CAL_DARM_DQ and H1:CAL-DELTAL_EXTERNAL_DQ. I calibrated these data with calibration files here. The results are shown in attachment 3. All the data and scripts are included in attachment 4, where analysis.py is the script that does the job. Based on the plots, it seems DAC noise could be potentially a limiting factor for the interferomter sensitivity.

The coil driver states for L1 is LP off, ACQ off (state 1). For H1 is LP on, ACQ off. The LISO files calculating the current transfer functions and the voltage transfer functions are attached in attachment 4. 

I used a resolution of 1mHz in the diaggui measurement. The data files are too large so I can not upload them here. I am figuring out what to do.

Note: I fell into a few traps during the calculation. Many of them was about data and transfer functions. I have been more careful about what data is used in these calculations. For example, the noisemon data downloaded from the sites when MASTER was off still has DAC noise in it. I thought it was ADC noise + noisemon noise before and used it for subtraction. Another example, the transfer function measured at the sites has all the noise in it. We do not see the noises in the passband but ADC noise dominates at high frequencies. If you use this transfer function to figure out how much noisemon noise contributes, you result will be tampered by the noises, like ADC noises at high frequency. Last example, if you use the noisemon noise data measured in the digital system in our lab, you should be aware that, although it does not have DAC noise (I disconnected DAC when measuring the noises), it also has ADC noise. Therefore, it would be better to use data from SR785 or LISO simulations (which has been shown to agree with each other). I drew a diagram in attachment 2 to help thinking about what data or transfer functions should be used. 

Quote:

you have to overlay the estimated displacemnt noise with the existing L1 noise bud or else we cant tell what the importance of the result is

 

Attachment 1: noisebudget.pdf
noisebudget.pdf
Attachment 2: diagram.pdf
diagram.pdf
Attachment 3: DACnoise-comparison.pdf
DACnoise-comparison.pdf
  1848   Wed Mar 11 12:46:20 2020 ranaComputingNoise BudgetNoisemon at L1

you have to overlay the estimated displacemnt noise with the existing L1 noise bud or else we cant tell what the importance of the result is

  1847   Thu Feb 13 00:25:36 2020 DuoDailyProgress Noisemon at L1

I calculated the DAC noise for L1. Attachment 1 has all the plots and data. Attachment 2 is the result in strain.

We have noisemon at all four stations: ITMX, ITMY, ETMX, ETMY. DTT gives me the CSD, the coherences and the ASD of all the channels at all the four stations. I use this to calculate the transfer functions of the coil driver and the noisemon.

H(f) = CSD_{ab}(f)/ D^2(f)

where D(f) is the drive and CSD is the CSD between the drive and the output of noisemon. The absolute value of H(f) will be the gain of the circuit, in ADC counts / DAC counts. Then I use the coherence to calculate the total noise

N_{drive}(f)=D(f)\sqrt{1 - C_{ab}(f)}

This noise has DAC noise, ADC noise, noisemon noise in it. This noise is in DAC counts. I will call this "DriveNoise".

Then I picked another time when the interferomter is not running and the drive is zero. I measured the noisemon spectrum N(f) at that time. The plots and data of these spectrum can be found in attachment 1. The plots are considered to be a result of the noisemon noise and ADC noise, which I will call "NoDriveNoise" (in ADC cts). Since the drive is zero, there is no DAC noise in it - just ADC noise and noisemon noise. I use the transfer function to covert it to DAC counts

N_{noDrive}(f)=N(f)/|H(f)|

Then I subtract the noise drive noise from the drive noise to get the DAC noise to get the DAC noise, which is then converted to DAC volts
N_{DAC}(f) = \sqrt{N_{drive}^2(f)-N_{noDrive}^2(f)}\frac{20V}{2^{18}cts}

Then I find the current on the coil using the transfer functions of the coil driver, assuming the coil driver is in LP OFF and ACQ OFF state. The transfer function can be found in attachment 1.

I(f)=N_{DAC}(f)*H_{I}(f)

where the transfer function H is a voltage-to-current transfer function.

Then we have the force

F(f)=I(f) \times 0.0309N/A

Lastly we have the displacement and strain

h(f)=\frac{F(f) (10Hz/f)^43\times10^{-8}}{4000}\sqrt{Hz}

For each station I summed all the four channels, LL, LR, UL, UR and then I calculated for all the four stations and summed as

h_{tot}(f)=\sqrt{h_{ITMX}^2(f)+h_{ITMY}^2(f)+h_{ETMX}^2(f)+h_{ETMY}^2(f))}

I tried to compare this with the GWINC model - it is much higher. I do not have real L1 noise at the moment. I will see once we have real noise data.

Attachment 1: L1DACNoise.zip
Attachment 2: plot.pdf
plot.pdf
  1846   Tue Feb 4 17:30:50 2020 DuoDailyProgressElectronicsSome confusing noisemon findings

I measured the spectrum with the digital system at all test points on noisemon with 10Hz sine drive (1845), but I saw a lot of distortion harmonics and other confusing stuff. At the end, I realized it could be caused by the clipper wire I used to breakout the DSUB9 connector. After talking to Chris today, I also realized that the low input impedance of the digital system could be invasive to the circuit. Considering these potential issues, I repeated yesterday's measurements with SR785 (attachment 3). There is not much difference - instead of using the digital system to send the drive and do the measurements, I used SR785 to drive and measure. SR785 has 1MOhm of input impedance while the ADC has only 10-20kOhm, according to Chris, so it could resolve the input impedance issue. Also, there is no DSUB connectors used; it also let me get rid of the clipper wires. 

I sent a 10Hz, 10mV p-p sine signal to the coil driver and measured all the test points from the output of the coil driver to the output of the noisemon. The results are in attachment 2, plotted in attachment 1. The test points can be referenced in the schematics in attachment 4. The plots with two numbers after 'TP' means the measurement is between the two test points - with postive clipper on one, negative on the other. Others are between the test point and ground. 

We can see there is a little bit distortion lines in the spectrum under 100Hz, but they are very small compared to the response to the drive. I think maybe they could explain the low frequency bump logged in 1843. However, it seems unlikely that they will be a serious issues in the passband, but still needs a little more rigorous justification.

Quote:

Trying to figure out the nature of the distortion found in noisemon (1843), I made a piece of DSUB9 breakout like attachment 1. I just used two wires and two clippers wires to break the DSUB connector into something I can clip into the board.

I sent a 10Hz drive to the coil driver and measured the spectrum at all the test points (including one pair on the coil driver board).

First, I compared the distortion situation at the input and output of noisemon (attachment 2). I measured the spectrum at TP10 and TP12 of coil driver (this is where noisemon picks up the differential input, reference DCC: D070483), and subtracted them quadratically, which gives us the input signal of noisemon. Then I measured the spectrum at the output of noisemon. Looking at attachment 2, it seems the coil driver does not contribute to the distortion.

Then I used the probe and measured all the test points individually. Attachment 3 also shows TP10 and TP12 but instead of measuring them individually and subtracting them, I just put one probe on TP10 and another on TP12. It looks like the coil driver does have a little distortion, contrast to the conclusion above.

Attachment 4 shows the spectrum at the noisemon input. There is a buffer between TP10/TP12 on coil driver and the input of noisemon (reference D070483). It looks like the distortion is less, somehow.

Attachment 5 is the spectrum after the passive filters, between TP3 and TP4. It seems the distortion comes back a little under 60Hz.

Attachment 6 shows the spectrum after the instrument amplifier.

Attachment 7 and 8 shows the output of two stages of high pass filters. The distortions gets much worse after two HP filters.

Attachment 9 shows the output of the low pass filter. We can see the high frequency harmonics are gone.

Attachment 10 is very confusing. Between TP8 and TP9 is just a buffer (LT1128 buffer unfortunately) - why does it give so many lines even without any drive? I tried to see them in the oscilloscope but there wasn't anything significant. Maybe it was the clippers/wires making the noise? I did check the signal at the conjunction point between the clipper and the wire using oscilloscope (attachment 11) - it is indeed bad (attachment 12) - but why these lines only show up in TP9?

I am very confused and need to think about what is going on now. I think a couple immediate questions are:

1. Why do I see so many lines at the output of noisemon even when there is no input? I should only see noisemon noise when there is no drive, but I did not short the inputs like when I measure the noise. I did shorted the inputs and checked the noise after the measurement - it is normal. Thus, the lines could be caused by not shorting the inputs of the coil driver, but I did not short the inputs in other measurements either - I just unchecked the excitation on diaggui.

2. Besides the TP9 confusion, we still see a lot of harmonics in uppersteam testpoints. Considering the TP9 situation, I think we should first ask - are they real? Then, if they are real, how bad are they? Considerations could include the functionality of noisemon board at the sites? This potentially include risk of saturation, increase of the noisemon noise and errors in measuring the DAC noise.

 

Attachment 1: p.pdf
p.pdf
Attachment 2: DataAndCode.zip
Attachment 3: Photo_on_2-4-20_at_5.27_PM.jpg
Photo_on_2-4-20_at_5.27_PM.jpg
Attachment 4: noisemon_v1.pdf
noisemon_v1.pdf noisemon_v1.pdf noisemon_v1.pdf noisemon_v1.pdf
  1845   Mon Feb 3 17:15:59 2020 DuoDailyProgressElectronicsSome confusing noisemon findings

Trying to figure out the nature of the distortion found in noisemon (1843), I made a piece of DSUB9 breakout like attachment 1. I just used two wires and two clippers wires to break the DSUB connector into something I can clip into the board.

I sent a 10Hz drive to the coil driver and measured the spectrum at all the test points (including one pair on the coil driver board).

First, I compared the distortion situation at the input and output of noisemon (attachment 2). I measured the spectrum at TP10 and TP12 of coil driver (this is where noisemon picks up the differential input, reference DCC: D070483), and subtracted them quadratically, which gives us the input signal of noisemon. Then I measured the spectrum at the output of noisemon. Looking at attachment 2, it seems the coil driver does not contribute to the distortion.

Then I used the probe and measured all the test points individually. Attachment 3 also shows TP10 and TP12 but instead of measuring them individually and subtracting them, I just put one probe on TP10 and another on TP12. It looks like the coil driver does have a little distortion, contrast to the conclusion above.

Attachment 4 shows the spectrum at the noisemon input. There is a buffer between TP10/TP12 on coil driver and the input of noisemon (reference D070483). It looks like the distortion is less, somehow.

Attachment 5 is the spectrum after the passive filters, between TP3 and TP4. It seems the distortion comes back a little under 60Hz.

Attachment 6 shows the spectrum after the instrument amplifier.

Attachment 7 and 8 shows the output of two stages of high pass filters. The distortions gets much worse after two HP filters.

Attachment 9 shows the output of the low pass filter. We can see the high frequency harmonics are gone.

Attachment 10 is very confusing. Between TP8 and TP9 is just a buffer (LT1128 buffer unfortunately) - why does it give so many lines even without any drive? I tried to see them in the oscilloscope but there wasn't anything significant. Maybe it was the clippers/wires making the noise? I did check the signal at the conjunction point between the clipper and the wire using oscilloscope (attachment 11) - it is indeed bad (attachment 12) - but why these lines only show up in TP9?

I am very confused and need to think about what is going on now. I think a couple immediate questions are:

1. Why do I see so many lines at the output of noisemon even when there is no input? I should only see noisemon noise when there is no drive, but I did not short the inputs like when I measure the noise. I did shorted the inputs and checked the noise after the measurement - it is normal. Thus, the lines could be caused by not shorting the inputs of the coil driver, but I did not short the inputs in other measurements either - I just unchecked the excitation on diaggui.

2. Besides the TP9 confusion, we still see a lot of harmonics in uppersteam testpoints. Considering the TP9 situation, I think we should first ask - are they real? Then, if they are real, how bad are they? Considerations could include the functionality of noisemon board at the sites? This potentially include risk of saturation, increase of the noisemon noise and errors in measuring the DAC noise.

Attachment 1: thumbnail_image0.jpg
thumbnail_image0.jpg
Attachment 2: distortionDiagnosis.pdf
distortionDiagnosis.pdf
Attachment 3: CoilDriverOutput.pdf
CoilDriverOutput.pdf
Attachment 4: NoisemonInput.pdf
NoisemonInput.pdf
Attachment 5: NoisemonPassiveFilter.pdf
NoisemonPassiveFilter.pdf
Attachment 6: NoisemonTP5.pdf
NoisemonTP5.pdf
Attachment 7: NoisemonTP6.pdf
NoisemonTP6.pdf
Attachment 8: NoisemonTP7.pdf
NoisemonTP7.pdf
Attachment 9: NoisemonTP8.pdf
NoisemonTP8.pdf
Attachment 10: NoisemonTP9.pdf
NoisemonTP9.pdf
Attachment 11: IMG_1685.jpg
IMG_1685.jpg
Attachment 12: IMG_1684.jpg
IMG_1684.jpg
  1844   Tue Jan 28 08:46:13 2020 ranaDailyProgressElectronicsDistortion of noisemon

it might be the low input impedance of the board which the coil driver cannot drive..

I suggest you use probes to see where in the noisemon circuit the distortion is starting

  1843   Tue Jan 21 16:13:03 2020 DuoDailyProgress Distortion of noisemon

We have sent version 2 noisemon boards (modifications from version 1 noted in 1835) to Livingston and Hanford. Chris noticed that there might be some upconversion problems under 20Hz (attachment 1 and 2). These plots from Chris are noisemon output with drive subtracted. Attachment 1 is the spectrum after replacement of noisemon board (version 2 board used), compared to attachment 2 which is before the replacement (version 1 board used). There is something going on near 15Hz. We think it might be due to upconversions under 15Hz.

We still have a spare board here. I modified it, tested it and looked at this issue, trying to reproduce. I sent a 10Hz 100 count amplitude signal to the board and compared it to the output with no input. This produces attachment 3. We see that once the 10Hz sine signal is sent, there are  lines above 10Hz, which is a sign of distortion.

I changed the input frequency to 5Hz and got attachment 4.

Attachment 5 is a linear plot to see where exactly those lines are. It seems they are indeed harmonics of 5Hz when the input signal is 5Hz.

I also tried higher frequencies up to 100Hz and saw similar harmonics.

 

Attachment 1: IMG_1658.PNG
IMG_1658.PNG
Attachment 2: IMG_1659.PNG
IMG_1659.PNG
Attachment 3: upconv.pdf
upconv.pdf
Attachment 4: upconv5Hz.pdf
upconv5Hz.pdf
Attachment 5: upconvLinear.pdf
upconvLinear.pdf
  1842   Tue Jan 14 17:07:49 2020 DuoDailyProgress Testing another two boards

I am testing another two boards to be shipped to Hanford today. I found and fixed some bad soldering. However, the frontend crashed and I needed to restart it. I could not log into cymac somehow. It says "no route to host" but the internet is still working. I cannot go on testing anymore. Let's wait and try again tomorrow.

Chris fixed the problem - I went on finishing the tests and the boards are ready to go now. They will be shipped tomorrow.

  1841   Mon Jan 13 16:48:52 2020 DuoDailyProgress Noisemon updates

I did some documentation work these days.

- Noisemon Test Plan: https://dcc.ligo.org/LIGO-E2000007

- Updated DCC, new schematics, PCB layout and BOM (due to changes logged in 1835)

- Two more boards are modified and will be tested.

  1840   Thu Dec 12 16:48:39 2019 DuoDailyProgress Modifying the board changes the transfer function

The noisemon board has been modified according to 1835. We do not expect any change in the transfer function but we see an increase in the gain above 20Hz.

Attachment 1 shows the comparison between modified board transfer function and the original board transfer function.

Attachment 3 shows the difference of the transfer functions in attachment 1. There is about 8-10dB increase at 100Hz after the modifications.

Attachment 4 is a comparison of LISO simulations. I calculated the transfer functions of the whole noisemon circuits before and after the modifications. I substracted them and found the difference.

Without putting attachment 3 and 4 in the same picture we can see that they are very different. LISO basically says there will not be any significant change in the transfer function but actually our measurement shows that there is.

Attachment 1: ModifiedVsIntact.pdf
ModifiedVsIntact.pdf
Attachment 2: ModifedVsIntact.xml.gz
Attachment 3: gainDiff.pdf
gainDiff.pdf
Attachment 4: lisocomparison.pdf
lisocomparison.pdf
  1839   Mon Dec 9 22:06:19 2019 DuoDailyProgress Two boards tested

I finished testing S1900294 and S1900297 and plan to ship them to the sites.

I got stuck at a couple things. I think it is good to make a note of these stuff.

1. Diaggui gives "unable to start excitation".

Solution: restart diaggui

2. Drive signal does not go in

Solution: check the status of the digital system - it crashed and needs restart.

3. FASTIMON gives too much noise that the transfer function looks like junk.

Solution: I use 50ohm resistors to replace the coils and 500 counts noise to measure the transfer function. If the resistance is large and the input signal is small, the current will be too small.

  1837   Thu Dec 5 20:42:55 2019 DuoDailyProgressElectronicsNoisemon board modifications

Fully board successful with modification. The saturation issue has been fixed on all the four channels.

Attachment 1: NoDrive.pdf
NoDrive.pdf
Attachment 2: 10HzSine.pdf
10HzSine.pdf
Attachment 3: TF.pdf
TF.pdf
  1836   Thu Dec 5 12:37:58 2019 Noisemon board modifiedDailyProgress Noisemon board modifications

The modification is successful. The attachment shows the result after the board runs for 12 hours. We modified channel 1 and 2 so we see FM0 and FM1 channels are still decent. We only split C2 but not R2 for channel 3 and 4 so FM2 and FM3 are bad.

Quote:

According to previous post 1834, we think the noisemon problem is very likely caused by C2 charging. Hence we did the following modifications on a noisemon board:

1. Split R2 into two 400 Ohm resistors, with grounding between them.

2. Split C2 into two 3.3uF capacitors, with grounding between them.

We hope these modifications will provide a path for the bias current to go to ground, instead of charging up C2.

 

Attachment 1: ModificationComparison.pdf
ModificationComparison.pdf
  1835   Wed Dec 4 19:00:37 2019 Noisemon board modifiedDailyProgress Noisemon board modifications

According to previous post 1834, we think the noisemon problem is very likely caused by C2 charging. Hence we did the following modifications on a noisemon board:

1. Split R2 into two 400 Ohm resistors, with grounding between them.

2. Split C2 into two 3.3uF capacitors, with grounding between them.

We hope these modifications will provide a path for the bias current to go to ground, instead of charging up C2.

  1834   Tue Dec 3 09:50:18 2019 DuoDailyProgress Interesting observation on noisemon debugging

We have observed that the noisemon bug happens after it is powered on for about 1.5 hour. Noisemon has been powered on overnight and the following morning I came in and found

- Channel 1, 3, 4 bad (with signals like attachment 1. Green: normal behavior driven with 250 count, 10Hz signal. Brown: abnormal behavior with no drive)

- Then I used the oscilliscope and did the following on channel 4:

1. Connected channel 1 to measure the voltage between one side of C2 and GND.

2. Connected channel 2 to measure the voltage between another side of C2 and GND.

3. Use MATH on the oscilliscope to measure the voltage difference

- Then I found channel 4 is good! I did not turn the power off or do anything else.

- I repeated exactly the same procedure on channel 3 and it is repeatable.

- I left channel 1 as comparison and made attachment 2.

- Then I just use the probe of oscilliscope to connect one side of C2 on channel 1 to GND and got attachment 3, which channel 1 is good again.

I think this is a very strong hint that this whole problem is due to C2 charging up.

Attachment 1: Saturating.pdf
Saturating.pdf
Attachment 2: diaggui.pdf
diaggui.pdf
Attachment 3: Goodagain.pdf
Goodagain.pdf
  1833   Mon Dec 2 15:28:03 2019 DuoMisc S1900296 board damaged

I think I shorted somewhere near the RTN testpoint on the board today while testing it. I saw some sparks. After that the board becomes non-responsive - it is not responding to whatever signal I send in. I will use another board and go on with testing.

  1832   Wed Nov 27 23:45:29 2019 DuoDailyProgress Noisemon Debug Progress

I found that two of the channels misbehaves after the board runs for a couple hours. Turning the power off and back on returns the circuit to normal functionality.

I sent 100 counts amplitude 10Hz sine into the board.

  • I saw channel 4 has some bad behavior - about 3V of offset with a lot of noise and one spike goes up to saturation.

Then I switched power off and back on. It worked normally. I increase amplitude to 2500, same as Carl's 10000. It worked normally as well. However, when I came back after a couple hours,

  • I saw two channels (1&4) saturates about 4 times a second. The other two still ok. (attachment 1)

 Then I turned power off and on again. I got attachment 2 - normal behavior again.

Attachment 1: saturating.pdf
saturating.pdf
Attachment 2: AfterRestart.pdf
AfterRestart.pdf
  1831   Mon Nov 25 16:43:50 2019 DuoDailyProgress Time domain projection is consistent with LLO data (except LL channel, which saturates)

We see a report of noisemon problem from LLO: https://alog.ligo-la.caltech.edu/aLOG/index.php?callRep=49892

The time domain data is projected with the transfer function measured here: https://alog.ligo-la.caltech.edu/aLOG/index.php?callRep=43212

We projected the output of the noisemon signal on time domain (attachment 1).

Attachment 2 is data from LLO posted in the alog above. Compare this with attachment 1, we can see our projection has roughly the same values with the other three channels. This means those channels have high number of counts because the drive signal is bigger. In other words, the other three channels should be working ok.

I attached the drive signal being projected in attachment 3.

Attachment 4 has all the data and code.

Attachment 1: timeresponse.pdf
timeresponse.pdf
Attachment 2: 49892_20191113153034_ITMY_NoiseMon.png
49892_20191113153034_ITMY_NoiseMon.png
Attachment 3: drive.pdf
drive.pdf
Attachment 4: archive.tar.gz
  1830   Thu Nov 14 17:14:32 2019 DuoDailyProgress Noisemon problem observation

Saturation occurs suddenly after running for hours without problem.

I tracked it since the drive was turned on (around 00:36:18) and found that it has been running ok for hours before this happens (attachment 1). Around 3:04:51, the LL channel saturated and then dropped back to normal level. After that, this kept happening and crashed the LL channel (attachment 3).

Attachment 2 is data corresponding to attachment 1.

Attachment 1: Screenshot_from_2019-11-14_17-13-46.png
Screenshot_from_2019-11-14_17-13-46.png
Attachment 2: ts.xml
<?xml version="1.0"?>
<!DOCTYPE LIGO_LW [
<!ELEMENT LIGO_LW ((LIGO_LW|Comment|Param|Time|Table|Array|Stream)*)>
<!ATTLIST LIGO_LW Name CDATA #IMPLIED Type CDATA #IMPLIED>
<!ELEMENT Comment (#PCDATA)>
<!ELEMENT Param (#PCDATA)>
<!ATTLIST Param Name CDATA #IMPLIED Type CDATA #IMPLIED Dim CDATA #IMPLIED
                Unit CDATA #IMPLIED>
<!ELEMENT Table (Comment?,Column*,Stream?)>
<!ATTLIST Table Name CDATA #IMPLIED Type CDATA #IMPLIED>
... 21160 more lines ...
Attachment 3: Screenshot_from_2019-11-14_17-26-55.png
Screenshot_from_2019-11-14_17-26-55.png
  1829   Thu Nov 14 15:54:57 2019 DuoDailyProgress Issue consistency among ETMY, ITMX and ITMY

I checked the behavior of noisemon in the three stations where they are installed. It is consistent that there is always one channels that is saturating and one channel that has excess noise. However, they are not always the same channels.

ETMY, attachment 1, LL saturating, UL excess noise (UR a little more noise)

ITMY, attachment 2, LL saturating, UL excess noise

ITMX, attachment 3, UL saturating, UR excess noise

Attachment 1: Screenshot_from_2019-11-14_15-54-27.png
Screenshot_from_2019-11-14_15-54-27.png
Attachment 2: Screenshot_from_2019-11-14_15-51-25.png
Screenshot_from_2019-11-14_15-51-25.png
Attachment 3: Screenshot_from_2019-11-14_15-51-15.png
Screenshot_from_2019-11-14_15-51-15.png
  1828   Thu Nov 14 15:04:25 2019 DuoDailyProgress ETMY check

The first noisemon board was installed at the ETMY station. It was a prototype board that we brought to LLO and installed there since then. I checked the data today (11/14) and its LL channels is not working.

I checked the time series in the attachment 2 and 3. There are some problem causing the circuit to saturate.

I checked the drive signal below 10Hz in attachment 4. The drive signals across all channels are the same.

Attachment 1: Screenshot_from_2019-11-14_15-02-05.png
Screenshot_from_2019-11-14_15-02-05.png
Attachment 2: Screenshot_from_2019-11-14_15-08-42.png
Screenshot_from_2019-11-14_15-08-42.png
Attachment 3: Screenshot_from_2019-11-14_15-15-56.png
Screenshot_from_2019-11-14_15-15-56.png
Attachment 4: Screenshot_from_2019-11-14_15-30-07.png
Screenshot_from_2019-11-14_15-30-07.png
  1827   Mon Nov 11 20:35:13 2019 DuoDailyProgress First sign of ITMY LL noisemon saturation

Noisemon has been installed in LLO ITMY station. Its LL channel starts to saturate at 10/30 00:00:00. I checked the noisemon output at that time (attachment 1).

Attachment 1: Screenshot_from_2019-11-11_20-33-46.png
Screenshot_from_2019-11-11_20-33-46.png
  1826   Fri Nov 8 13:32:15 2019 DuoDailyProgress Noisemon and FastIMON at L1

I checked the October 25 data at L1. The drive signal is zero, according to the bottom plot in attachment 1 and 2. The fastimon channel has a lot of lines. Noisemon still has the 65Hz harmonics.

I also checked the Nov 7 noisemon and fastimon data. We still see the harmonics on Nov 7. Also, one channel is completely dysfunctioning. The fastimon is similar to Oct 25.

Attachment 1: Oct 25 noisemon. We still see the harmonics in the noise.

Attachment 2: Oct 25 fastimon. Two channels are messy.

Attachment 3,4,5: Nov 7 noisemon data. Analyzed in attachment 4. Data in attachment 5.

Attachment 6: Nov 7 fastimon data. Similar to Oct 25 - a lot of lines.

Attachment 1: Screenshot_from_2019-11-08_13-14-42.png
Screenshot_from_2019-11-08_13-14-42.png
Attachment 2: Screenshot_from_2019-11-08_13-11-51.png
Screenshot_from_2019-11-08_13-11-51.png
Attachment 3: Screenshot_from_2019-11-08_13-53-45.png
Screenshot_from_2019-11-08_13-53-45.png
Attachment 4: nm1107.pdf
nm1107.pdf
Attachment 5: Desktop.zip
Attachment 6: Screenshot_from_2019-11-08_14-22-20.png
Screenshot_from_2019-11-08_14-22-20.png
  1825   Wed Oct 23 20:37:53 2019 DuoDailyProgress Nosiemon check

Although L1 has been down since the beginning of October, I checked the spectrum at the output of the nosiemon at L1 ITMY. It looks like the UL channel, as we have seen before, still has the lines,

Attachment 1: Screenshot_from_2019-10-23_20-48-13.png
Screenshot_from_2019-10-23_20-48-13.png
  1824   Tue Oct 8 17:31:27 2019 DuoSummary Noisemon check

Noisemon has been installed at L1 for a while. Now we have it on ITMX and ITMY. ITMX was installed first and ITMY was installed on September 11.

Recently, after it was installed at ITMY, we were trying to check the functionality of the circuit - if it measures DAC nosie properly between 20 - 100Hz. When we were doing that, we encountered some strange harmonics of 65Hz in UL channel. It is shown in attachment 6, data from 9/25 ETMY.

We traced back to where it was installed at ETMY, 9/11, and we still see the lines there (attachment 5).

Then we went to check ITMX, since the data was consistent when it was first installed. However, when we use 9/25 data, we see attachment 4. The UL channel measurement is way above the other three channels. Tracing back the dates, we found in ITMX it started 9/10 (attachment 3), comparing to 9/9 (attachment 2) when it was still good.

All the data and scripts are in attachment 1.

Notes (in case you read about noisemon for the first time):

All the plots are

drive\times \sqrt{1-coherence}

It is suppose to be DAC nosie + Noisemon noise + ADC noise, which DAC noise being dominant between 20 - 100Hz. The purple curves are DAC noise model. The measurements are expected to be close to it between 20Hz and 100Hz.

 

Attachment 1: Desktop.zip
Attachment 2: itmx99.pdf
itmx99.pdf
Attachment 3: itmx910.pdf
itmx910.pdf
Attachment 4: itmx925.pdf
itmx925.pdf
Attachment 5: itmy911.pdf
itmy911.pdf
Attachment 6: itmy925.pdf
itmy925.pdf
  1823   Wed Oct 2 13:50:03 2019 DuoDailyProgress ITMY noisemon

This is the DAC noise of ITMY PUM coil driver in L1. The data is from September 25, 2019 00:00:00 UTC. I did not subtract the ADC noise and the noisemon noise since I did it for ITMX. The noisemon noise is far below the total noise, and the ADC noise only dominates high frequency. After subtraction, the mismatch above 500Hz will disappear.

There are some lines around 65Hz. I checked the data a few days arounds 9/25. I see them in all the data. I do not understand why yet.

Spectra of the PUM drive signal is attached in 3, noisemon output in 4. I do not see anything at 65Hz in those channels.

ADC noise and Noisemon noise are subtracted in attachment 5.

I checked these data

9/23 00:00, 8:00, 20:00

9/24 8:00,

9/25 0000,0800,2000

9/26 0800,2000

9/27 0000.0800.2000

10/1 0800

The 65Hz harmonics stays there.

I checked time domain for 9/15. It is not saturating. (Attachment 6&7)

Attachment 8 is a check on the output of the noisemon signal when L1 is down. It is from 9/24 1600. Surprisingly, I see the 65Hz harmonics in UL spectrum while I cannot see it when the interferometer is online. I do not see the harmonics in the other three channels.

In attachment 9, I checked ITMX from the same time (9/24 1600, interferometer down). The 65Hz harmonics is there in all channels, and the UL output is higher.

Attachment 1: plot.pdf
plot.pdf
Attachment 2: Desktop.zip
Attachment 3: drive.pdf
drive.pdf
Attachment 4: nm.pdf
nm.pdf
Attachment 5: plot.pdf
plot.pdf
Attachment 6: Archive.zip
Attachment 7: ts.png
ts.png
Attachment 8: Screenshot_from_2019-10-04_16-43-25.png
Screenshot_from_2019-10-04_16-43-25.png
Attachment 9: Screenshot_from_2019-10-04_17-11-12.png
Screenshot_from_2019-10-04_17-11-12.png
  1822   Wed Sep 11 17:24:43 2019 DuoSummaryElectronicsITMX DAC noise

I did a quick estimation of the subtraction result. I subtracted the ADC noise and the noisemon noise. 

NoSubtraction=\sqrt{1-coherence}\times drive

Now I subtract it:

WithSubtractionDACNoise=\sqrt{NoSubtraction^2-NoisemonNoise^2-ADCNoise^2}

The ADC noise and Noisemon noise are converted to DAC volts (divided by the transfer function of the noisemon and coil driver).

Form the results from 10-200Hz, it seems that the calculated noise is DAC noise.

Attachment 2 is the result: subtracted noises vs. model compared to aLIGO noise.

Attachment 1 shows how the subtraction is done: we subtract noisemon noise from the total noise. (Noisemon noise contains ADC noise) Noisemon noise and ADC noise is neglegible at that frequency.

It seems at low frequency what we see there might still be DAC noise, if not other unknown sources.

Quote:

I don't agree about this. Doesn;t this ignore the noise of the noisemon circuit (analog readout noise + ADC noise) ? I think you must have a model for than noise in order to infer the DAC noise. Or maybe my pringle suggestion has better SNR?

Quote:

This is how we calculate the DAC noise spectrum. The unit is V/rtHz.

drive\times\sqrt{1-coherence}

 

 

 

Attachment 1: plot.pdf
plot.pdf
Attachment 2: beforeSub.pdf
beforeSub.pdf
  1821   Sat Aug 31 19:21:13 2019 ranaSummaryElectronicsITMX DAC noise

I don't agree about this. Doesn;t this ignore the noise of the noisemon circuit (analog readout noise + ADC noise) ? I think you must have a model for than noise in order to infer the DAC noise. Or maybe my pringle suggestion has better SNR?

Quote:

This is how we calculate the DAC noise spectrum. The unit is V/rtHz.

drive\times\sqrt{1-coherence}

 

 

  1820   Fri Aug 30 14:45:42 2019 ranaComputingDAQDownload data with pyNDS

I'm attaching a script to download data from the LIGO sites with python.

I recommend using it in your anaconda3 ENV:

conda install -c conda-forge nds2-client python-nds2-client

and then before running the script you have to initialize your Kerberos token:

kinit miley.cyrus@LIGO.ORG

then you run the script:

python getData.py --ifo=L1 --fs=1024

as usual, run with the -O or -OO flags to silence the debug messages.

Attachment 1: ChanList_darm.txt
SUS-ETMY_L2_MASTER_OUT_LL_DQ
SUS-ETMY_L2_NOISEMON_LL_OUT_DQ
Attachment 2: getData.py
#!/usr/bin/env python
# this function gets some data (from the 40m) and saves it as
# a .mat file for the matlabs
# Ex. python -O getData.py


import scipy.io as sio
import scipy.signal as sig
from astropy.time import Time
import nds2
... 101 more lines ...
  1819   Fri Aug 30 13:15:07 2019 DuoSummaryElectronicsITMX DAC noise

The PUM noisemon board has been installed in Livingston ITMX test mass. After the installation, we fetched the coil driver drive signal, noisemon output signal and the coherence between them. 

This is how we calculate the DAC noise spectrum. The unit is V/rtHz.

drive\times\sqrt{1-coherence}

The data fetching configurations are:

- Start time: 1250467218. Locked for more than 20hrs from there, if you check here: https://ldas-jobs.ligo-la.caltech.edu/~detchar/summary/day/20190822/

- Bin size: 0.1Hz

- Window: Hanning

- Average: 1000

Attachment 1 and 2: plot of the DAC noises in volts compared to the G1401399 model.

Attachment 3 and 4: plot of the DAC noises projected to the displacement of the test mass and incohrerently summed from all the four test masses.

All the data are attached as xml files. They are directly saved from DTT and can be opened in DTT.

To reproduce the plots, run the python code in the zip file. The code runs without any parameters. 

Attachment 1: plot.pdf
plot.pdf
Attachment 2: Desktop.zip
Attachment 3: projected.pdf
projected.pdf
Attachment 4: Archive.zip
ELOG V3.1.3-